# Tidal disruption of a neutron star near naked singularity

### v2

@@ V Summary
The potential for bright kilonova-like transients to be triggered from NS disruption in supermassive compact objects without horizons is a novel aspect of this study. Depending on how much of the disrupted mass escapes, such events could yield significant r-process material and possibly enrich the universe with more heavy elements than typical NS–NS mergers. The observational signatures of these events are significantly shaped by the relativistic timescales for both bound and unbound debris, which we also discuss. Crucially, the two scenarios have different expected light curves: for BHs, it rises to a maximum and then decays, whereas for NaSs, because of the lack of an event horizon, it keeps diverging.
@@ VI Acknowledgment
Appendix A Geometrical structure of static Spacetimes
@@
A.1 Matching condition for finite boundary spacetime
@@
A.2 Joshi-Malafarina-Narayan 1 NaS
@@
Appendix B Jacobi fields on timelike radial geodesics in static spherically symmetric (SSS) spacetimes
@@
B.1 Tidal force and Jacobi field in the JMN1 NaS spacetime

### v3

@@ V Summary
The potential for bright kilonova-like transients to be triggered from NS disruption in supermassive compact objects without horizons is a novel aspect of this study. Depending on how much of the disrupted mass escapes, such events could yield significant r-process material and possibly enrich the universe with more heavy elements than typical NS–NS mergers. The observational signatures of these events are significantly shaped by the relativistic timescales for both bound and unbound debris, which we also discussed. Crucially, the two scenarios have different expected light curves: for BHs, it rises to a maximum and then decays, whereas for NaSs, because of the lack of an event horizon, it keeps diverging.
@@ Acknowledgements.
Appendix B Geometrical structure of static Spacetimes
@@
B.1 Matching condition for finite boundary spacetime
@@
B.2 Joshi-Malafarina-Narayan type 1 Naked Singularity
@@
Appendix C Jacobi fields on timelike radial geodesics in static spherically symmetric (SSS) spacetimes
@@
C.1 Tidal force and Jacobi field in the JMN1 Naked Singularity spacetime